OpenStudy (anonymous):

can you factorize it for me please (l+m)2 - (l-m)2

OpenStudy (bahrom7893):

This is basically a^2-b^2 = (a-b)(a+b) so: (l+m)2 - (l-m)2 = [(l+m) + (l-m)][(l+m) - (l-m)]

OpenStudy (anonymous):

or 2lm = (l+m)^2 - (l-m)^2

OpenStudy (bahrom7893):

now just simplify.. [(l+m) + (l-m)][(l+m) - (l-m)] = [2l][2m] = 4lm

OpenStudy (anonymous):

(l+m)2 - (l-m)2 = [(l+m) + (l-m)][(l+m) - (l-m)] = [2l] [2m] = 4 *l*m

OpenStudy (anonymous):

\[a^2-b^2=(a+b)(a-b)\] use \[a=l+m,b=l-m\]

OpenStudy (anonymous):

4lm = (l+m)^2 - (l-m)^2

OpenStudy (anonymous):

when you substitute you get \[(l+m+l-m)(l+m-(l-m)\] \[=(2l)(2m)=4lm\] just like everyone else said